Ten times the first number exceeds the second by five.
Represent the numbers by n and 7n + 19.
Find the numbers.
(I dont know how to set up the equation to solve)

Answer by ankor@dixie-net.com(22740)

Ten times the first number exceeds the second by five.
Represent the numbers by n and 7n + 19.
Find the numbers.
:
n = the 1st number
and
(7n+19) = the 2nd number
:
The equation
10n = (7n+19) + 5
10n = 7n + 24
10n - 7n = 24
3n = 24
n = 8
:
:
See if that is true in the original equation
10(8) = (7(8) + 19) + 5
80 = 56 + 19 + 5